OpenStudy (justuu):

Medal & Fan Which function has a y-value of 6, if the x -value is 3? y = x - 3 y = 2x y = x + 2 y = 3x

OpenStudy (dtan5457):

From the last lesson i'm sure you can answer this, no?

Nnesha (nnesha):

again replace x by 3

Nnesha (nnesha):

then solve after solving which equation give you 6 that is your answer

OpenStudy (dtan5457):

y = (3) - 3 y = 2(3) y = (3) + 2 y = 3(3)

OpenStudy (dtan5457):

Which of those equal 6?

OpenStudy (justuu):

okay i got it

Nnesha (nnesha):

which one is your answer ??

Nnesha (nnesha):

just for make sure :D

OpenStudy (dtan5457):

to make sure ^

OpenStudy (dtan5457):

yes :P

Nnesha (nnesha):

let us know if you have question :) :)